# Optimal Clustering with Dependent Costs in Bayesian Networks

### v4

@@ Appendix A Supporting Information: Super-additive Objective Functions
Appendix B Supporting Information: Simple Example
@@
Note that all three optimal solutions place nodes $D$ and $G$ in the same cluster (cluster 2, shaded black in Fig. 1). $D$ is not only the node with the highest in-degree, but also connects the top part of the graph with the middle and, via node $G$, the bottom part of the graph. Thus it is feasible for $D$ and $G$ to be in the same cluster. Nodes $B$, $C$ and $E$ (shades of gray in Fig. 1) are all in individual clusters, corresponding to root ($B$, $C$) nodes or a node along a linear chain ($E$). As the objective function is based on computation cost, it is expected that root nodes are best assigned to their own clusters (Lemma A.1). The key variation between the three least-cost cluster mappings is the mapping of node $A$ which can either be part of cluster 2 with $D$ and $G$, part of cluster 1 with node $F$, or in an independent cluster of its own. These solutions were discovered at iterations 35, 182, and 276 on branches 21, 27 and 104, respectively. With a cost of 54.0, this is much less than the naive heuristic solution cost of 85.8.
@@
Although the earlier solutions are quite different to the optimal solution, the solution at iteration 21 starts to follow the trend of the optimal solution with $D$, $G$ and $A$ in the same cluster, with other nodes mostly in their own clusters. This suggests that finding near-optimal or even the first optimal solution could potentially be very rapid. As might be expected, however, finding all optimal solutions could be expensive (Likhachev et al., 2005; Belfodil et al., 2019). In addition, near-optimal solutions could produce diverse cluster mappings, although they converge towards optimal mapping patterns fairly rapidly.

### v5

@@ 6 Supplementary Information: Super-additive Objective Functions
7 Supplementary Information: Simple Example
@@
Note that all three optimal solutions place nodes $D$ and $G$ in the same cluster (cluster 2, shaded black in Fig. 1). $D$ is not only the node with the highest in-degree, but also connects the top part of the graph with the middle and, via node $G$, the bottom part of the graph. Thus it is feasible for $D$ and $G$ to be in the same cluster. Nodes $B$, $C$ and $E$ (shades of gray in Fig. 1) are all in individual clusters, corresponding to root ($B$, $C$) nodes or a node along a linear chain ($E$). As the objective function is based on computation cost, it is expected that root nodes are best assigned to their own clusters (Lemma 6.1). The key variation between the three least-cost cluster mappings is the mapping of node $A$ which can either be part of cluster 2 with $D$ and $G$, part of cluster 1 with node $F$, or in an independent cluster of its own. These solutions were discovered at iterations 35, 182, and 276 on branches 21, 27 and 104, respectively. With a cost of 54.0, this is much less than the naive heuristic solution cost of 85.8.
@@
Although the earlier solutions are quite different to the optimal solution, the solution at iteration 21 starts to follow the trend of the optimal solution with $D$, $G$ and $A$ in the same cluster, with other nodes mostly in their own clusters. This suggests that finding near-optimal or even the first optimal solution could potentially be very rapid. As might be expected, however, finding all optimal solutions could be expensive [17, 2]. In addition, near-optimal solutions could produce diverse cluster mappings, although they converge towards optimal mapping patterns fairly rapidly.
Declarations
  • Competing Interests: There are no competing interests.